# Integral Cayley graphs over a finite symmetric algebra

### v1

@@ 1. Introduction
In [6], Klotz and Sander study unitary Cayley graphs over $\operatorname{\mathbb{Z}}/n$ and their natural generalization: the gcd-graphs $G_{n}(D)$ where $D$ is a subset of proper divisors of $n$. Let us quickly recall the definition of $G_{n}(D)$. The vertices of $G_{n}(D)$ are elements of the finite ring $\operatorname{\mathbb{Z}}/n$ and two elements $a,b$ are adjacent if $\gcd(a-b,n)\in D.$ Using the theory of Ramanujan sums, Klotz and Sander show that the gcd-graphs $G_{n}(D)$ are integral. They also ask whether the converse is also true; namely, if a Cayley graph over $\operatorname{\mathbb{Z}}/n$ is integral, it is true that it is a gcd-graph? In [11], So provides an affirmative answer to this question. In [9], inspired by the analogy between number fields and function fields, we study gcd-graphs over polynomial rings. Among various things that we find, we show that the spectrum of these gcd-graphs also has an explicit formula via Ramanujan sums. An important insight of our work is the notion of a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra (see [9, Definition 6.2]). By their very definition, symmetric algebras are self-dual, and consequently, their characters are parametrized by their elements, which allows us to calculate the associated Gauss and Ramanujan sums quite explicitly.
@@
We remark that our main theorem works for all $S$ but since we only work with undirected graphs, we impose the above conditions on $S$. When $R=\operatorname{\mathbb{Z}}/n$, this recovers So’s theorem. We note, however, that our theorem applies to a much wider class of finite rings. More precisely, as we will show in the last section, all finite quotients of the ring of integers in a global field are symmetric. We believe that this observation could pave the way for future studies on gcd-graphs over such rings as well as Paley graphs associated with finite Hecke characters.
@@ 2. Main results
Let $R$ be a finite symmetric $\operatorname{\mathbb{Z}}/n$-algebra with a fixed non-degenerate linear functional $\psi.$ Then for each every character $\widehat{\psi}$ of $R$, there exists a unique element $r\in R$ such that for all $t\in R$
@@
By definition, $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{d}$ are linear independent over $\operatorname{\mathbb{Q}}$. Let $\mathcal{V}$ be the $\operatorname{\mathbb{Q}}$-vector space generated by the $\bm{v}_{i}$’s.
@@
By 2.4 we know that $A_{R}\mathcal{A}\subset\mathcal{V}.$ Since $A_{R}$ is invertible, we conclude that $\dim(\mathcal{A})\leq\dim(\mathcal{V}).$ Since $\mathcal{V}\subset\mathcal{A}$, we conclude that $\mathcal{A}=\mathcal{V}.$
@@
The forward direction has been proved in 2.1. Let us prove the other direction. Suppose that $\Gamma(R,S)$ is integral. Then by 2.5, $\bm{1}_{S}\in\mathcal{V}$. Since the components of $\bm{1}_{S}$ are in the set $\{0,1\}$, we conclude that $S$ is a union of some of the orbits $O_{1},O_{2},\ldots,O_{d}$. In other words, $S$ is stable under the action of $(\operatorname{\mathbb{Z}}/n)^{\times}.$
@@
We then see that 2.6 is a generalization of [11, Theorem 6.2].
@@ 3. Examples of symmetric algebras
In this section, we provide some constructions of finite symmetric $\operatorname{\mathbb{Z}}/n$-algebras. The first example is quite standard (see [7, Example 3.15E]).
@@
Since $\operatorname{\mathcal{O}}_{K}$ is a DVR, $\mathcal{I}$ is a principal ideal; namely $\mathcal{I}=\langle f\rangle$ for some $f\in\operatorname{\mathcal{O}}_{K}.$ We then have, $\operatorname{\mathcal{O}}_{K}/\mathcal{I}\cong\operatorname{\mathcal{O}}_{K}/f.$ By 3.5, it is enough to show that $\operatorname{\mathcal{O}}_{K}/p^{a}$ is a symmetric $\operatorname{\mathbb{Z}}/p^{a}$-algebra. It is well-known that $\operatorname{\mathcal{O}}_{K}$ is monogenic over $\operatorname{\mathbb{Z}}_{p}$, i.e, there is $\alpha\in\operatorname{\mathcal{O}}_{K}$ such that $\operatorname{\mathcal{O}}_{K}=\operatorname{\mathbb{Z}}_{p}[\alpha]$ (see [10, Chapter 3, Section 6, Proposition 12]). Let $m=[K:\operatorname{\mathbb{Q}}_{p}]$, then every element in $\operatorname{\mathcal{O}}_{K}/p^{a}$ can be written uniquely as $\sum_{i=0}^{m-1}a_{i}\alpha^{i}$ where $\alpha_{i}\in\operatorname{\mathbb{Z}}_{p}/p^{a}=\operatorname{\mathbb{Z}}/p^{a}.$ We can define a linear functional $\psi\colon\operatorname{\mathcal{O}}_{K}/p^{a}\to\operatorname{\mathbb{Z}}/p^{a}$ by
@@
By an identical argument for the proof of [9, Proposition 6.7], we can check that $\psi$ is non-degenerate and hence $\operatorname{\mathcal{O}}_{K}/p^{a}$ is a symmetric $\operatorname{\mathbb{Z}}/p^{a}$-algebra. ∎
@@
A local field of characteristic $p$ is isomorphic to the Laurent series $\operatorname{\mathbb{F}}_{q}((t))$ for some $q$. As a result, $\operatorname{\mathcal{O}}_{K}\cong\operatorname{\mathbb{F}}_{q}[[t]]$. We can then find $a$ such that $\mathcal{I}=\langle t^{a}\rangle.$ We then have $\operatorname{\mathbb{F}}_{q}[[t]]/\mathcal{I}\cong\operatorname{\mathbb{F}}_{q}[[t]]/t^{a}\cong\operatorname{\mathbb{F}}_{q}[t]/t^{a}.$ By [9, Corollary 6.8], $\operatorname{\mathbb{F}}_{q}[t]/t^{a}$ is a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra. ∎
@@
By 2.6 and 3.9, we have the following corollary which answers a question posed in [9, Remark 6.22].
@@ 3.1. Integral generalized Paley graphs
Let $\operatorname{\mathcal{O}}_{K}$ be the ring of integers of a number field $K.$ Let $\mathcal{I}$ be a non-zero ideal of $\operatorname{\mathcal{O}}_{K}$. Suppose $\chi\colon(\operatorname{\mathcal{O}}_{K}/\mathcal{I})^{\times}\to\operatorname{\mathbb{C}}^{\times}$ is a character such that $\chi(-1)=1$. The following definition is motivated by the definition of the quadratic Paley graphs as defined and studied in [8] and [2, Section 4.2].
@@
There are various constructions of $\chi$ with the property that the induced Dirichlet character on $(\operatorname{\mathbb{Z}}/n)^{\times}$ is trivial. For example, let us consider $K=\operatorname{\mathbb{Z}}[i]$. For each $n\in\operatorname{\mathbb{Z}}$, let $\chi_{n}:(\operatorname{\mathbb{Z}}[i]/n)^{\times}\to\operatorname{\mathbb{C}}^{\times}$ be the quartic residue symbol associated with $n$ (see [4, Chatper 9]). Then, by [4, Proposition 9.8.3], we know that $\chi_{n}(a)=1$ for all $a\in(\operatorname{\mathbb{Z}}/n)^{\times}.$ As a result, the Paley graph $P_{\chi}$ is integral.
It seems to be interesting to study some further arithmetic properties of the eigenvalues of $P_{\chi}$ in this case. We hope to do so in a future work.
@@ Acknowledgements
We thank Minh-Tam Trinh for introducing us to symmetric algebras and for some clarifications on Hecke characters. We also thank Ján Mináč for his constant encouragement and support.

### v2

@@ 1. Introduction
In [8], Klotz and Sander study unitary Cayley graphs over $\operatorname{\mathbb{Z}}/n$ and their natural generalization: the gcd-graphs $G_{n}(D)$ where $D$ is a subset of proper divisors of $n$. Let us quickly recall the definition of $G_{n}(D)$. The vertices of $G_{n}(D)$ are elements of the finite ring $\operatorname{\mathbb{Z}}/n$ and two elements $a,b$ are adjacent if $\gcd(a-b,n)\in D.$ Using the theory of Ramanujan sums, Klotz and Sander show that the gcd-graphs $G_{n}(D)$ are integral. They also ask whether the converse is also true; namely, if a Cayley graph over $\operatorname{\mathbb{Z}}/n$ is integral, is it true that it is a gcd-graph? So (2016) in [13] provides an affirmative answer to this question. For the reader’s convenience, we recall So’s theorem ([13, Theorem 7.1]) in terms of Cayley graphs over $\operatorname{\mathbb{Z}}/n$ as follows.
Theorem (So’s theorem).
A undirected Cayley graph $G=\Gamma(\operatorname{\mathbb{Z}}/n,S(G))$ is integral if and only if $S(G)$ is a union of the $G_{n}(d)^{\prime}s$.
Here for each $d\mid n$,
$G_{n}(d)=\{m\in\operatorname{\mathbb{Z}}/n\mid\gcd(m,n)=d\}.$
In [11], inspired by the analogy between number fields and function fields, we study gcd-graphs over polynomial rings. Among various things that we find, we show that the spectrum of these gcd-graphs also has an explicit formula via Ramanujan sums. An important insight of our work is the notion of a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra (see [11, Definition 6.2]). By their very definition, symmetric algebras are self-dual, and consequently, their characters are parametrized by their elements, which allows us to calculate the associated Gauss and Ramanujan sums quite explicitly.
@@
We remark that our main theorem works for all $S$ but since we only work with undirected graphs, we impose the above conditions on $S$. When $R=\operatorname{\mathbb{Z}}/n$, this recovers So’s theorem. We note, however, that our theorem applies to a much wider class of finite rings. More precisely, as we will show in the last section, all finite quotients of the ring of integers in a global field are symmetric. We believe that this observation could pave the way for future studies on gcd-graphs over such rings as well as Paley graphs associated with finite Hecke characters (see 3.13).
@@ 2. Main results
Example 2.3.
$\operatorname{\mathbb{Z}}/n$ is a symmetric $\operatorname{\mathbb{Z}}/n$-algebra, where the identity map from $\operatorname{\mathbb{Z}}/n$ to itself is a non-degenerate linear functional. More generally, as we will show in Section 3, a finite quotient of the ring of integers in a global field is a symmetric algebra.
@@
Let $R$ be a finite symmetric $\operatorname{\mathbb{Z}}/n$-algebra with a fixed non-degenerate linear functional $\psi.$ Then for each character $\widehat{\psi}$ of $R$, there exists a unique element $r\in R$ such that for all $t\in R$
@@
By definition, $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{d}$ are linearly independent over $\operatorname{\mathbb{Q}}$. Let $\mathcal{V}$ be the $\operatorname{\mathbb{Q}}$-vector space generated by the $\bm{v}_{i}$’s.
@@
By 2.5 we know that $A_{R}\mathcal{A}\subset\mathcal{V}.$ Since $A_{R}$ is invertible, we conclude that $\dim(\mathcal{A})\leq\dim(\mathcal{V}).$ Since $\mathcal{V}\subset\mathcal{A}$, we conclude that $\mathcal{A}=\mathcal{V}.$
@@
The forward direction has been proved in 2.1. Let us prove the other direction. Suppose that $\Gamma(R,S)$ is integral. Then by 2.6, $\bm{1}_{S}\in\mathcal{V}$. Since the components of $\bm{1}_{S}$ are in the set $\{0,1\}$, we conclude that $S$ is a union of some of the orbits $O_{1},O_{2},\ldots,O_{d}$. In other words, $S$ is stable under the action of $(\operatorname{\mathbb{Z}}/n)^{\times}.$
@@
We then see that 2.7 is a generalization of [13, Theorem 7.1].
@@ 3. Examples of symmetric algebras
In this section, we provide some constructions of finite symmetric $\operatorname{\mathbb{Z}}/n$-algebras. The first example is quite standard (see [9, Example 3.15E]).
@@
Since $\operatorname{\mathcal{O}}_{K}$ is a DVR, $\mathcal{I}$ is a principal ideal; namely $\mathcal{I}=\langle f\rangle$ for some $f\in\operatorname{\mathcal{O}}_{K}.$ We then have, $\operatorname{\mathcal{O}}_{K}/\mathcal{I}\cong\operatorname{\mathcal{O}}_{K}/f.$ By 3.5, it is enough to show that $\operatorname{\mathcal{O}}_{K}/p^{a}$ is a symmetric $\operatorname{\mathbb{Z}}/p^{a}$-algebra. It is well-known that $\operatorname{\mathcal{O}}_{K}$ is monogenic over $\operatorname{\mathbb{Z}}_{p}$, i.e, there is $\alpha\in\operatorname{\mathcal{O}}_{K}$ such that $\operatorname{\mathcal{O}}_{K}=\operatorname{\mathbb{Z}}_{p}[\alpha]$ (see [12, Chapter 3, Section 6, Proposition 12]). Let $m=[K:\operatorname{\mathbb{Q}}_{p}]$, then every element in $\operatorname{\mathcal{O}}_{K}/p^{a}$ can be written uniquely as $\sum_{i=0}^{m-1}a_{i}\alpha^{i}$ where $\alpha_{i}\in\operatorname{\mathbb{Z}}_{p}/p^{a}=\operatorname{\mathbb{Z}}/p^{a}.$ We can define a linear functional $\psi\colon\operatorname{\mathcal{O}}_{K}/p^{a}\to\operatorname{\mathbb{Z}}/p^{a}$ by
@@
By an identical argument for the proof of [11, Proposition 6.7], we can check that $\psi$ is non-degenerate and hence $\operatorname{\mathcal{O}}_{K}/p^{a}$ is a symmetric $\operatorname{\mathbb{Z}}/p^{a}$-algebra. ∎
@@
A local field of characteristic $p$ is isomorphic to the Laurent series $\operatorname{\mathbb{F}}_{q}((t))$ for some $q$. As a result, $\operatorname{\mathcal{O}}_{K}\cong\operatorname{\mathbb{F}}_{q}[[t]]$. We can then find $a$ such that $\mathcal{I}=\langle t^{a}\rangle.$ We then have $\operatorname{\mathbb{F}}_{q}[[t]]/\mathcal{I}\cong\operatorname{\mathbb{F}}_{q}[[t]]/t^{a}\cong\operatorname{\mathbb{F}}_{q}[t]/t^{a}.$ By [11, Corollary 6.8], $\operatorname{\mathbb{F}}_{q}[t]/t^{a}$ is a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra. ∎
@@
By 2.7 and 3.9, we have the following corollary which answers a question posed in [11, Remark 6.22].
@@ 3.1. Integral generalized Paley graphs
Let $\operatorname{\mathcal{O}}_{K}$ be the ring of integers of a number field $K.$ Let $\mathcal{I}$ be a non-zero ideal of $\operatorname{\mathcal{O}}_{K}$. Suppose $\chi\colon(\operatorname{\mathcal{O}}_{K}/\mathcal{I})^{\times}\to\operatorname{\mathbb{C}}^{\times}$ is a character such that $\chi(-1)=1$. The following definition is motivated by the definition of the quadratic Paley graphs as defined and studied in [10] and [3, Section 4.2].
@@
There are various constructions of $\chi$ with the property that the induced Dirichlet character on $(\operatorname{\mathbb{Z}}/n)^{\times}$ is trivial. For example, let us consider $K=\operatorname{\mathbb{Z}}[i]$. For each $n\in\operatorname{\mathbb{Z}}$, let $\chi_{n}:(\operatorname{\mathbb{Z}}[i]/n)^{\times}\to\operatorname{\mathbb{C}}^{\times}$ be the quartic residue symbol associated with $n$ (see [6, Chatper 9]). Then, by [6, Proposition 9.8.3], we know that $\chi_{n}(a)=1$ for all $a\in(\operatorname{\mathbb{Z}}/n)^{\times}.$ As a result, the Paley graph $P_{\chi}$ is integral. It seems to be interesting to study some further arithmetic properties of the eigenvalues of $P_{\chi}$ in this case. For example, one may ask whether it is possible to calculate the spectrum of $P_{\chi}$ explicitly (the case in which $\chi$ is a quadratic Dirichlet character has been solved in [10].) Furthermore, since $P_{\chi}$ is an integral graph, it is a promising candidate for the existence of perfect state transfers (see [2, 4]). We, therefore, wonder whether it is possible to classify $P_{\chi}$ which admit perfect state transfers. We hope to discuss these questions in future work.
@@ Acknowledgements
We thank Minh-Tam Trinh for introducing us to symmetric algebras and for some clarifications on Hecke characters. We also thank Ján Mináč for his constant encouragement and support. Finally, we thank the anonymous referee for their help in improving the quality and clarity of this paper.